OpenStudy (dottiecat):

What is the slope of the line passing through the points (-1, 3) and (4, -7)?\ A 2 B -2 C -4/3 D 3/4

OpenStudy (dottiecat):

Hi again

OpenStudy (mecharv):

Remember the previous graph? Same concept. The two points are given here... so you can substitute in the same formula

OpenStudy (dottiecat):

What should I do with the negatives

OpenStudy (dottiecat):

-1 - 4 / 3 - -7

OpenStudy (mecharv):

-1-4=-5 3-(-7)=3+7=10 Two negatives make a positive :)

OpenStudy (dottiecat):

Oh yeah :D

OpenStudy (dottiecat):

so then it's 5/10. divide 5 by 10. then it's 0.5...?

OpenStudy (mecharv):

No, you have to take the reciprocal and check whether it is 5 or -5? O_o

OpenStudy (dottiecat):

oh

OpenStudy (mecharv):

m=10/-5=-2

OpenStudy (dottiecat):

ohh, lol i was working it out on paper. but thanks anyways XD

OpenStudy (mecharv):

Anytime :)